# HoloNet: Toward a Unified Einstein–Maxwell–Dilaton Framework of QCD

### v1

@@ Abstract
We propose HoloNet, a neural-network framework that unifies lattice QCD(LQCD) thermodynamics and holographic Einstein–Maxwell–Dilaton (EMD) theory within a data-to-holography pipeline. Instead of assuming specific functional forms, HoloNet learns the metric profile $A(z)$ and the gauge–dilaton coupling $f(\phi)$ directly from 2+1-flavor LQCD data at $\mu=0$. These learned functions are embedded into the EMD equations, enabling the model to reproduce the lattice equation of state and baryon number fluctuations with high fidelity. Once trained, HoloNet provides a fully data-driven holographic description of QCD that extends naturally to finite density, allowing us to map the phase diagram and estimate the location of the critical end point (CEP). The reconstructed potential $V(\phi)$ and coupling $f(\phi)$ agree quantitatively with those obtained from holographic renormalization, demonstrating that HoloNet can consistently bridge different holographic models.
@@ I Introduction
At the same time, the rapid development of machine learning aarts2025physics has endowed holographic models with significantly enhanced representational capabilities. In particular, the neural network approach enables a more efficient and effective solution of ordinary differential equations, especially for models in which the equations of motion contain functional dependencies that are difficult to handle analytically or numerically using traditional methods song2021ads; ahn2025holographic; chen2018neural; chen2025data; hashimoto2025machine; hashimoto2024unification; ahn2024deep. By introducing neural networks along the holographic direction Hashimoto:2020jug; akutagawa2020deep; hashimoto2021neural, it becomes possible to use boundary data and reconstruct the bulk metric. Using this approach, we can construct holographic models driven directly by LQCD data.
@@ II HOLOGRAPHIC EMD MODEL
$\mathcal{L}$ denotes the Lagrangian density.
When the metric functions $A(z)$ and coupling functions $f(\phi)$ are undetermined, the equations of motion can be formally expressed in the following integral form. Based on Ref. Yang:2014bqa, we have simplified the relevant expressions to make them more suitable for constructing the neural network,
@@ III Neural-Network Approach
Figure 2: The graphs of the AA and $f$ functions are shown, with red representing the optimized neural network results, black corresponding to an analytically guessed solution, blue depicting the machine learning results obtained from a prescribed functional form Chen:2024ckb, and green indicating the outcomes of Bayesian analysis zhu2025bayesian. Additionally, the gray areas in the figures represent the $z$ range over which the lattice data are available.
@@
Meanwhile, empirically, the function $A(z)$ should be a monotonically decreasing function. Therefore, we use the softplus activation function in the last layer, which helps the neural network architecture as much as possible conform to this prior behavior.
@@
Figure 3: The figures display the speed of sound and specific heat, with red solid lines representing our analytical results and the error bars indicating the lattice data bazavov2014equation. As anticipated, these quantities exhibit excellent agreement once the model accurately satisfies the EoS.
@@ V Reconstruction Results
The LQCD data are taken from  bazavov2014equation; borsanyi2012fluctuations. From the LQCD data, through the optimization of the neural networks, we can reconstruct the bulk functions $A(z)$ and $f(z)$, as shown in Fig. 2.
@@ V.1 Thermodynamic Observables
Fig. 4 compares the LQCD data (error bars) at zero chemical potential with the optimized neural network results (solid lines). The comparison covers four quantities: energy density, entropy density, pressure, and anomaly. Among these, the entropy density is used to optimize the network, while the energy density, pressure and anomaly automatically match the lattice results after optimization. At zero chemical potential, all these quantities depend only on the metric function $A(z)$.
Figure 4: EoS comparison between neural network results and LQCD data bazavov2014equation. The error bars represent the LQCD data, while the solid lines represent the neural network model results. The blue curve corresponds to the trace anomaly $\epsilon-3p$, the yellow to the entropy density $s$, the green to the pressure $p$, and the red to the energy density $\epsilon$. The entropy density and pressure are used in training the neural network model. Since there are only two independent thermodynamic variables, the trace anomaly and energy density generally agree with the lattice data automatically.
@@
Figure 5: Baryon number susceptibility comparison between neural network result and LQCD data borsanyi2012fluctuations. The black dots represent the LQCD results, with error bars omitted due to their negligible size. The red solid line represents the neural network model results. The optimized model exhibits a high level of quantitative agreement with the lattice results.
@@ V.2 Closed-Form Expression
Figure 6: Reconstructed potential function $V(\phi)$ from neural network training. The black dashed line represents the potential obtained from the holographic renormalization method, while the blue solid line represents the potential obtained from the potential reconstruction neural network model. Within the $\phi$ range used in the numerical calculations, the values of $V(\phi)$ show a high degree of agreement.
Figure 7: Coupling function $f(\phi)$ from neural network training. The black dashed line represents the coupling function obtained from the holographic renormalization method, while the blue solid line represents the coupling function obtained from the potential reconstruction neural network model. Within the $\phi$ range involved in the numerical calculations, the values of $f(\phi)$ are in close agreement.
We can reconstruct the potential functions $V(\phi)$ and $f(\phi)$ using the learned $A(z)$ and $f(z)$. The results are shown in Fig. 6 and Fig. 7. The reconstructed potential functions exhibit quantitative agreement with those obtained through the holographic renormalization method Cai:2022omk. It should be noted that in the $\phi\to 0$ region, the holographic renormalization scheme introduces a pole in $f(\phi)$, leading to a slight difference between the two approaches. This pole, in  Cai:2022omk, was artificially introduced for computational convenience, has no physical significance, and can be removed. This consistency serves as a crucial validation of our neural network approach. Eventually, our reconstructions support the following analytical forms of $V(\phi)$) and $f(\phi)$,
@@ V.3 Extrapolation to CEP
Figure 8: The phase diagram is shown, where the red solid and dashed lines denote the first-order phase transition and crossover obtained from the neural network, and the black line indicates the analytical guess. The gray-shaded region represents the range in which the model predictions are reliable, obtained by extending the zero-chemical-potential equation of state to finite chemical potential through the baryon number susceptibility in the holographic model. The light-blue region indicates the part of the phase diagram where the CEP has been excluded borsanyi2025lattice. Green ones represent DeWolfe-Gubser-Rosen-type HQCD models, and green left triangle grefa2021hot, upward triangle zhao2023phase, and downward triangle hippert2024bayesian correspond to related HQCD results. Blue dot, solid, and dashed lines denote the CEP, first-order lines, and crossovers from chen2025flavor. The pink and yellow regions show the $68\%$ and $95\%$ confidence-level CEP estimates from zhu2025bayesian, with the blue pentagram indicating the maximum a posteriori result. yellow right triangle, purple square, and purple diamond represent the CEPs obtained from the PNJL li2019kurtosis, FRG fu2020qcd, and DSE-FRG gao2020qcd approaches, respectively.
@@ Acknowledgements
We thank Dr. Daichi Takeda and Dr. Zhibin Li for helpful discussions. We thank the DEEP-IN working group at RIKEN-iTHEMS for support in the preparation of this paper. LW is supported by the RIKEN-TRIP initiative (RIKEN Quantum), JSPS KAKENHI Grant No. 25H01560, and JST-BOOST Grant No.JPMJBY24H9; MH is supported in part by the National Natural Science Foundation of China (NSFC) Grant Nos: 12235016, 12221005. HZ is supported by the Science and Technology Development Plan Project of Jilin Province, China Grant No. 20240101326JC.

### v2

# HoloNet: Toward a Unified Einstein-Maxwell-Dilaton Framework of QCD

@@ Abstract
We propose HoloNet, a neural-network framework that unifies lattice QCD(LQCD) thermodynamics and holographic Einstein-Maxwell-Dilaton (EMD) theory within a data-to-holography pipeline. Instead of assuming specific functional forms, HoloNet learns the metric profile $A(z)$ and the gauge–dilaton coupling $f(z)$ directly from 2+1-flavor LQCD data at $\mu=0$. These learned functions are embedded into the EMD equations, enabling the model to reproduce the lattice equation of state and baryon number fluctuations with high fidelity. Once trained, HoloNet provides a fully data-driven holographic description of QCD that extends naturally to finite density, allowing us to map the phase diagram and estimate the location of the critical end point (CEP). The reconstructed potential $V(\phi)$ and coupling $f(\phi)$ agree quantitatively with those obtained from holographic renormalization, demonstrating that HoloNet can consistently bridge different holographic models.
@@ I Introduction
At the same time, the rapid development of machine learning aarts2025physics has endowed holographic models with significantly enhanced representational capabilities. In particular, the neural network approach enables a more efficient and effective solution of ordinary differential equations, especially for models in which the equations of motion contain functional dependencies that are difficult to handle analytically or numerically using traditional methods song2021ads; jeong2025ads; ahn2025holographic; chen2018neural; chen2025data; hashimoto2025machine; hashimoto2024unification; ahn2024deep; ahn2025deep. By introducing neural networks along the holographic direction Hashimoto:2020jug; akutagawa2020deep; hashimoto2021neural, it becomes possible to use boundary data and reconstruct the bulk metric. Using this approach, we can construct holographic models driven directly by LQCD data.
@@ II HOLOGRAPHIC EMD MODEL
$\mathcal{L}$ denotes the Lagrangian density. When the metric functions $A(z)$ and coupling functions $f(\phi)$ are undetermined, the equations of motion can be formally expressed in the following integral form. Based on Ref. Yang:2014bqa, we have simplified the relevant expressions to make them more suitable for constructing the neural network,
@@ III Neural-Network Approach
Figure 2: The graphs of the AA and $f$ functions are shown, with red representing the optimized HoloNet results, black corresponding to an analytically guessed solution, blue depicting the machine learning results obtained from a prescribed functional form Chen:2024ckb, and green indicating the outcomes of Bayesian analysis zhu2025bayesian. Additionally, the gray areas in the figures represent the $z$ range over which the lattice data are available.
@@
Meanwhile, empirically, the function $A(z)$ should be a monotonically decreasing function zhang2022spectra. Therefore, we use the softplus activation function in the last layer, which helps the neural network architecture as much as possible conform to this prior behavior.
@@
Figure 3: The figures display the speed of sound and specific heat, with red solid lines representing our HoloNet results and the error bars indicating the lattice data bazavov2014equation. As anticipated, these quantities exhibit excellent agreement once the model accurately satisfies the EoS.
@@ V Reconstruction Results
The $(2+1)$-flavor LQCD data are taken from  bazavov2014equation; borsanyi2012fluctuations. From the LQCD data, through the optimization of the neural networks, we can reconstruct the bulk functions $A(z)$ and $f(z)$, as shown in Fig. 2.
@@ V.1 Thermodynamic Observables
Fig. 4 compares the LQCD data (error bars) at zero chemical potential with the optimized HoloNet results (solid lines). The comparison covers four quantities: energy density, entropy density, pressure, and anomaly. Among these, the entropy density is used to optimize the network, while the energy density, pressure, and anomaly automatically match the lattice results after optimization. At zero chemical potential, all these quantities depend only on the metric function $A(z)$.
Figure 4: EoS comparison between neural network results and LQCD data bazavov2014equation. The error bars represent the LQCD data, while the solid lines represent the HoloNet model results. The blue curve corresponds to the trace anomaly $\epsilon-3p$, the yellow to the entropy density $s$, the green to the pressure $p$, and the red to the energy density $\epsilon$. The entropy density is used in training the HoloNet model. Since these are non-independent thermodynamic variables, the pressure, trace anomaly, and energy density generally agree with the lattice data automatically.
@@
Figure 5: Baryon number susceptibility comparison between HoloNet result and LQCD data borsanyi2012fluctuations. The black dots represent the LQCD results, with error bars omitted due to their negligible size. The red solid line represents the neural network model results. The optimized model exhibits a high level of quantitative agreement with the lattice results.
@@ V.2 Closed-Form Expression
Figure 6: Reconstructed potential function $V(\phi)$ from HoloNet training(red). By contrast, the black line is an analytical closed-form result. The yellow dashed line represents the potential obtained from the holographic renormalization method, while the blue and green dashed lines represent the results from chen2025flavor and zhu2025bayesian, respectively. Within the $\phi$ range used in the numerical calculations, the values of $V(\phi)$ show a high degree of agreement.
Figure 7: Coupling function $f(\phi)$ from neural network training(red). The black line is an analytical closed-form result. The yellow solid line represents the coupling function obtained from the holographic renormalization method, while the blue and green dashed lines represent the results from chen2025flavor and zhu2025bayesian. Within the $\phi$ range involved in the numerical calculations, the values of $f(\phi)$ are in close agreement with the holographic renormalization method.
We can reconstruct the potential functions $V(\phi)$ and $f(\phi)$ using the learned $A(z)$ and $f(z)$. The results are shown in Fig. 6 and Fig. 7. The reconstructed potential functions exhibit quantitative agreement with those obtained through the holographic renormalization method Cai:2022omk. It should be noted that in the $\phi\to 0$ region, the holographic renormalization scheme introduces a pole in $f(\phi)$, leading to a slight difference between the two approaches. This pole, in  Cai:2022omk, was artificially introduced for computational convenience by making $f(0)=1$, has no physical significance, and can be removed. This consistency serves as a crucial validation of our neural network approach. Eventually, our reconstructions support the following analytical closed forms of $V(\phi)$) and $f(\phi)$,
@@ V.3 Extrapolation to CEP
Figure 8: The phase diagram is shown, where the red solid and dashed lines denote the first-order phase transition and crossover obtained from the neural network, and the black line indicates the analytical guess. The gray-shaded region represents the range in which the model predictions are reliable, obtained by extending the zero-chemical-potential equation of state to finite chemical potential through the baryon number susceptibility in the holographic model. The light-blue region indicates the part of the phase diagram where the CEP has been excluded borsanyi2025lattice. Green ones represent DeWolfe-Gubser-Rosen-type HQCD models, and green left triangle grefa2021hot, upward triangle zhao2023phase, and downward triangle hippert2024bayesian correspond to related HQCD results. Blue dot and solid line denote the CEP and first-order lines from chen2025flavor. The orange and light orange regions show the $68\%$ and $95\%$ confidence-level CEP estimates from zhu2025bayesian, with the dark orange pentagram indicating the maximum a posteriori result. The yellow right triangle is the result of PNJL li2019kurtosis. Purple square, and purple diamond represent the CEPs obtained from the FRG fu2020qcd, and DSE-FRG gao2020qcd approaches, respectively. The results of the further DSE–FRG approach are shown as orange hexagons gao2021chiral and purple crosses gunkel2021locating.
@@ Acknowledgements
We thank Drs. Daichi Takeda, Zhibin Li, and Jan M. Pawlowski for helpful discussions. We thank the DEEP-IN working group at RIKEN-iTHEMS for support in the preparation of this paper. LW is supported by the RIKEN-TRIP initiative (RIKEN Quantum), JSPS KAKENHI Grant No. 25H01560, and JST-BOOST Grant No.JPMJBY24H9; MH is supported in part by the National Natural Science Foundation of China (NSFC) Grant Nos: 12235016, 12221005. HZ is supported by the Science and Technology Development Plan Project of Jilin Province, China Grant No. 20240101326JC.